OpenStudy (anonymous):

6/x+4 +4/x+4=-5 what is the solution set

jhonyy9 (jhonyy9):

us parantheses please

jhonyy9 (jhonyy9):

if x+4 is denominator than you get 6+4=-5(x+4) 10=-5x-20 30=-5x x=30/(-5) x= -6